OpenStudy (anonymous):

9(x+2) > 9(x-3) true ?

OpenStudy (anonymous):

@Brainybeauty

OpenStudy (anonymous):

Solve it using the distributive property and you'll be able to tell

OpenStudy (anonymous):

When u solve the left side, 9(x+2) = 9x + 18 When u solve the right side, 9(x-3) = 9x - 27

OpenStudy (anonymous):

is the left side 9 and the right side 36 ?

OpenStudy (anonymous):

Adding 27 to both sides, Left Side = 9x + 45 Right Side = 9x 9x + 45 is always greater than 9x => Left Side > Right Side => 9(x+2) > 9(x-3) So it is true